OpenStudy (anonymous):

(1+kx)(64-192x-240x^2)=0 Expand and Find the value of k in which in there is no term in x^2

OpenStudy (anonymous):

find the product

OpenStudy (anonymous):

the sum of the coefficients of x^2 is 0

OpenStudy (anonymous):

the productt of what sorry i dont understand

OpenStudy (anonymous):

product of (1+kx) and (64-192x -240x^2)

OpenStudy (anonymous):

ok, u have not to find the whole answer . u know that u will get x^2, by multiplying -240x^2 and 1 , -192x and kx

OpenStudy (anonymous):

so,-240 -192k=0

OpenStudy (campbell_st):

look at the terms in x^2 so its 1*(-240x^2) and kx*(-192x) to find the value of k... equate the terms and solve for k \[-240x^2 = 192kx^2\] k = -240/192 \[k =- \frac{5}{4}\]

OpenStudy (anonymous):

cheers :)